OpenStudy (anonymous):

(x+6)(x−6)=x^​2+cx+d Find the values for integers c and d that makes the following equation true.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (xapproachesinfinity):

foil the left side and equate

OpenStudy (anonymous):

my b didnt see u

OpenStudy (anonymous):

soo how would tht go im still trying to get familiar with it

OpenStudy (xapproachesinfinity):

(x+6)(x-6)=x^2-36

OpenStudy (xapproachesinfinity):

you know that (a+b)(a-b)=a^2-b^2 yes ?

OpenStudy (anonymous):

wait where u get 32 from

OpenStudy (xapproachesinfinity):

difference of squares do you remember that formula?

OpenStudy (anonymous):

uhhhhhhhhhhh maybe:/

OpenStudy (xapproachesinfinity):

no may be yes or no

OpenStudy (anonymous):

nope

OpenStudy (xapproachesinfinity):

ok let's go slow then can you distribute (a+b)(a-b)

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

ok go ahead

OpenStudy (anonymous):

a^2-b^2

OpenStudy (xapproachesinfinity):

ok what about (x+6)(x-6)

OpenStudy (anonymous):

x^2-16

OpenStudy (xapproachesinfinity):

where did 16 come from?

OpenStudy (anonymous):

6+6 ohhh 6^2 soo x^2-6^2

OpenStudy (xapproachesinfinity):

does 6+6=6^2?

OpenStudy (anonymous):

no brain fart my b

OpenStudy (xapproachesinfinity):

and 6+6=12 not 16?

OpenStudy (xapproachesinfinity):

what is 6x6

OpenStudy (anonymous):

36

OpenStudy (xapproachesinfinity):

ok! do you know that 6^2 is the same as 6x6

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

then why did you tell me 6+6=6^2?

OpenStudy (anonymous):

didnt mean to i ment to say x^2-6^2

OpenStudy (xapproachesinfinity):

ok so x^-6^2=...

OpenStudy (anonymous):

u mean \[x ^{2}-6^{2}\]

OpenStudy (xapproachesinfinity):

yes

OpenStudy (xapproachesinfinity):

so \[x^2-6^2=x^2-36\] yes?

OpenStudy (anonymous):

(x+6)(x−6)

OpenStudy (anonymous):

damn back at beginning

OpenStudy (xapproachesinfinity):

yes that's what we started with \[(x+6)(x-6)=x^2-6^2=x^2-36\]

OpenStudy (anonymous):

ok thats better now i know where 36 came from

OpenStudy (xapproachesinfinity):

so we have now \[x^2-36=x^2+cx+d \]

OpenStudy (xapproachesinfinity):

do you know what coefficients mean?

OpenStudy (anonymous):

2 pairs right

OpenStudy (xapproachesinfinity):

no

OpenStudy (xapproachesinfinity):

why are doing this if you don't any idea of what a coefficient mean? did you not study this ?

OpenStudy (anonymous):

dang A number or symbol multiplied with a variable or an unknown quantity in an algebraic term.

OpenStudy (xapproachesinfinity):

ok yes that is a coef now tell me the coefficients of this expression \[x^2+cx+d\]

OpenStudy (anonymous):

x?

OpenStudy (xapproachesinfinity):

well x is the variable or unknown it can't be a coefficient

OpenStudy (anonymous):

idk

OpenStudy (xapproachesinfinity):

the coefficient for this expression are 1, c and d

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

read that definition of a coefficient and see if this is true a coefficient is a number or symbol that is multiplied to unknown 1 is multilpied to x^2 c is multiplied to x d is multiplied to x^0 (hidden unknown )

OpenStudy (xapproachesinfinity):

ok if you understand the concept what are the coefficients of this one x^2-36

OpenStudy (anonymous):

yes its true

OpenStudy (anonymous):

uhhhh x,2,36

OpenStudy (xapproachesinfinity):

incorrect

OpenStudy (xapproachesinfinity):

x is unknown 2 is a power not a coeff 36 is good but you missed - sign

OpenStudy (anonymous):

hmmmk

OpenStudy (xapproachesinfinity):

go watch some tutorial videos about this you need to wok on this before you try this problem

OpenStudy (anonymous):

ight

OpenStudy (xapproachesinfinity):

i could give you the answer but no point in that if you don't got why it is the correct one first get the basics and then go to this level of problem

OpenStudy (xapproachesinfinity):

https://www.youtube.com/watch?v=T9vz755kd20